# Where to Go Next Day: Multi-scale Spatial-Temporal Decoupled Model for Mid-term Human Mobility Prediction

### v1

@@ IV-D Trajectory Decoder
The trajectory decoder is comprised of a spatial decoder and a temporal decoder. Considering the dependency between duration and location, we expect to predict the location chain first and then utilize it to enhance the prediction of the duration chain. Inspired by the natural language process tasks, we add the start and end tokens into the location chain to address the challenges of the unknown start location and the uncertain chain length. Besides, we employ the Transformer decoder [48] as the backbone of the spatial decoder, along with a positional encoding module and an FNN. The self-attention mechanism in the Transformer decoder and positional encodings enable the model to effectively capture long-range dependencies and maintain the temporal order of the input sequence, thereby alleviating accumulative errors in multi-step prediction tasks. Commencing from the start token $\rm<SOS>$, the spatial decoder utilizes the historical information and predicted results to make iterative predictions for the daily location chain. This process is formulated as:
@@
$\hat{\rm C}^{\rm dur}=\{\hat{t}_{1}\dots,\hat{t}_{m}\}={\rm TemporalDec}(\{\hat{s^{\prime}}_{1}\oplus z^{\rm dur},\dots,\hat{s^{\prime}}_{m}\oplus z^{\rm dur}\}),$ (10)
@@ V-A1 Data Description
Experiments are conducted on four private  [^1] The dataset is not publicly disclosed due to privacy constraints. urban-scale datasets, Boston, Los Angeles(LA), San Francisco Bay Area(SFBay), Shanghai, collected from real-world mobile records, and one public dataset Tokyo [49], comprising synthetic mobility data. The preprocessing steps for the CDR datasets follow [24]. Besides, we partition each city into uniform grids with grid sizes of 1 kilometer. As described in Preliminary, a day is divided into multiple time windows, and only the record with the longest stay duration is retained within each window. Boston, LA, and SFBay utilize half-hour windows, whereas Shanghai and Tokyo use hour-long windows. The statistics of the datasets are presented in Table I, where “# Days” means the average number of days and “# Admins” means the number of the administrative division. Boston, LA, and SFBay employ census tracts as administrative units, whereas Shanghai uses towns. In the Tokyo dataset, due to blurred exact coordinates, we partitioned administrative regions using 7x7 grid squares. The training, validation, and test sets are obtained by partitioning the original dataset in a ratio of 6:1:3 based on the time span.
@@
Datasets # Individuals # Locations # Admins # Days # Records
Boston 50,000 10,294 1,126 39 6,749,540
LA 50,000 6,188 2,493 41 7,768,373
SFBay 50,000 8,032 1,406 41 7,700,986
Shanghai 20,000 16,050 234 99 8,209,910
Tokyo 16,151 7,153 225 75 7,502,839
@@ V-A2 Baselines
Seven models for predicting next locations and two models for successive mobility prediction are utilized for comparative analysis. They are CFPRec [15], GETNext [16], GFlash [50], HGARN [51], SNPM [41], AGRAN [42], MTNet [33], HTAED [38], and RLMob [39]. A concise introduction to these baselines is listed as follows:
@@
  • HTAED [38] introduces an LSTM-based encoder-decoder model with a hierarchical temporal attention module for successive mobility prediction.
@@ V-A4 Implementation
These baselines are implemented with their open-source codes. Hyperparameter tuning is performed on the validation set to attain optimal performance. The implementation of MSTDP is based on the PyTorch framework. The heterogeneous GNNs adopt a two-layer structure, while the layer numbers of Transformer encoder and decoder are also set to 2. The embedding dimensions of location and time are 512 and 512. The hidden dimensions of location and time are 1024 and 512. The head numbers of attention in the Transformer encoder and decoder are set to 8. We apply the Adam optimizer to train MSTDP with a learning rate of $10^{-4}$. The hyperparameter $\lambda$ is set to 1.
@@ V-B Metric Comparison
We initially evaluate all models using datasets Boston, Shanghai, and Tokyo, and subsequently select the more recent baselines to compare against datasets LA and SFBay. The results are exhibited in Table II and Table III. From these results, we draw the following observations and analysis: (1) MSTDP demonstrates superior performance on both datasets, validating its effectiveness in mid-term prediction tasks. Compared to the top-performing baseline, MSTDP exhibits improvements of 3.6% in Acc, 9.0% in DevDist, 20.3% in TravelDist and 19.4% in DepartTime. A smaller DevDist of MSTDP indicates closer proximity between the predicted and the actual locations. Furthermore, lower metrics for TravelDist and DepartTime demonstrate improved individual consistency between predicted and actual trajectories regarding travel time and distance patterns in our model. (2) Among these baselines, HTAED excels due to its dual-layer attention mechanism tailored for successive prediction, effectively weighing intra-day and inter-day behavioral patterns. In contrast, RLMob, which employs a reinforcement learning framework for successive prediction, falls short by relying solely on RNNs to model transition relationships, resulting in less favorable performances. These baselines, initially developed for next-location prediction, overlook temporal patterns such as daily and weekly cycles, which complicates their direct application to mid-term forecasting tasks. Moreover, when applied to Boston, LA, or SFBay datasets, these baseline predictions show significant performance degradation compared to Shanghai and Tokyo, notably highlighted by SNPM. This discrepancy primarily stems from varying time window sizes across datasets; for instance, Boston employs a half-hour window, leading to increased location redundancy and longer daily trajectories. These factors present substantial challenges for model performance, as discussed in Introduction I.
@@ V-C Statistical Physics Analysis
Origin-Destination (OD) flow pairs. We collect the cell-level OD flow pairs between the actual and predicted flows and illustrate them in Figure 7. We calculate $\rm{R}^{2}$ and CPC [52] between actual and predicted flows for numerical comparison. AGRAN and MTNet predict a limited number of flows, consistent with previous analyses indicating their preference for predicting full-day stays. HTAED overestimates flows with an $\rm R^{2}$ of -0.442 and $\rm CPC$ of 0.461, whereas our MSTDP model achieves closer predictions to actual flow quantities, with an improved $\rm R^{2}$ of 0.728 and $\rm CPC$ of 0.672.
@@
Origin-Destination flows in map. Figure 8 illustrates the actual and predicted OD flows in the map, where the trips between cells are aggregated at the census tract level. The global travel patterns predicted by the three baseline models deviate significantly from the actual results. HTAED achieves $\rm R^{2}$ and CPC values of 0.534 and 0.629, whereas MSTDP achieves 0.872 and 0.792, indicating performance improvements of 63.3% and 25.9% in our model. In this subsection, we examine OD flows at the cell and tract levels. Additionally, Figure 9 illustrates OD flows at the zip code level. MSTDP achieves R-squared ($R^{2}$) and CPC values of 0.969 and 0.877, while HTAED obtains 0.928 and 0.817. The performance improvements are 4.4% and 7.3%, respectively. This observation suggests that larger spatial aggregations may obscure differences between models.
@@ V-D Ablation Study
We compare MSTDP with three variants to examine the effectiveness of the modules in MSTDP. The ablated variants of MSTDP are as follows: (1) MSTDP-D: it eliminates the spatial-temporal decoupler and directly encodes the long trajectories; (2) MSTDP-H: it abandons the hierarchical structure and solely utilizes the daily encoder; (3) MSTDP-G: it replaces the graph-based location representation with a word embedder; (4) MSTDP-LD: the temporal decoder excludes the concatenation from location decoder and uses a Transformer decoder structure rather than the Transformer encoder.
@@
Next Day Prediction MSTDP-D 0.571 3.891 0.539 0.560
@@
Next Week Prediction MSTDP-D 0.404 5.634 0.564 0.570
@@ V-E Hyperparameter Analysis
To examine the MSTDP’s sensitivity to hyperparameters, we vary two hyperparameters to observe performance fluctuation. $\lambda$ adjusts the weight of the temporal loss, impacting the optimization process. We test $\lambda$ at values of 0.01, 0.1, 1, 10, and 100. Additionally, we also evaluate the head number (HN) of the attention in the Transformer encoder, considering values of 2, 4, and 8.
Table V presents the performance fluctuation with the hyperparameter variations. The results indicate that increasing the number of attention heads improves model performance. The influence of $\lambda$ on performance fluctuations is relatively modest; larger values such as 1 or 100 yield superior results compared to smaller values like 0.1 or 0.01. This highlights the critical role of duration time learning.
@@
Next Day Prediction #$\#Head=2$ 0.568 3.826 0.392 0.426
#$\#Head=4$ 0.571 3.803 0.389 0.430
@@
$\lambda=1,\#Head=8$ 0.584 3.676 0.361 0.419
Next Week Prediction #$\#Head=2$ 0.405 5.441 0.470 0.443
#$\#Head=4$ 0.403 5.478 0.468 0.450
$\lambda=0.01$ 0.404 5.470 0.467 0.445
$\lambda=0.1$ 0.411 5.349 0.469 0.447
$\lambda=10$ 0.407 5.445 0.465 0.444
$\lambda=100$ 0.418 5.292 0.473 0.436
$\lambda=1,\#Head=8$ 0.413 5.369 0.455 0.438
@@ V-F Application to Epidemic Transmission
The SEIR model is commonly used to represent the epidemic spreading in a population [53, 22]. It categorizes all individuals into four states: susceptible (S), exposed (E), infectious (I), and removed (R). Susceptible individuals can become infected through close contact with those carrying the infection and subsequently transition into the exposed state. There exists a probability of transitioning from the exposed state to infection. Lastly, infectious individuals transition into the removed state with a probability. The removed state assumes that individuals are no longer infected and cannot transmit the infection to others. Specifically, the probability of being exposed is formulated as $\rm P_{exp,i}=\frac{\alpha\beta I_{i}}{\rm N_{i}}$, where $\alpha$ is the number of close contacts, $\beta$ is the infectious rate, $\rm I_{i}$ and $\rm N_{i}$ stand for the number of infectious people and total population in the $i$-th subregion of the city. The infectious probability $\rm P_{inf}$ is the inverse of the average incubation period, while the removed probability $\rm P_{rem}$ is the inverse of the infectious period. The detailed parameter settings are illustrated in Table VI. $\rm R_{0}$, the basic reproduction number, represents the average number of secondary infections caused by an infectious individual. It is calculated as the product of $\alpha$,$\beta$, and the average infectious period.
@@
In this study, due to the sparse population in the dataset, we initially aggregate cell-level trajectories to the admin level (census tract for Boston) and assign an SEIR epidemic model to each administrative area. Subsequently, we employ these trajectories to construct a population transition probability matrix for each time window. These matrices are normalized to represent the probability of individuals choosing to stay or move between administrative areas during different periods. The population of each administrative area is obtained from Worldpop  [^2] https://www.worldpop.org/. For each time window, we first employ the SEIR model to calculate the status change of individuals within each administrative area, and then determine individual movement based on the transition probability matrix. All individuals are considered independent of each other.

### v2

@@ IV-D Trajectory Decoder
The trajectory decoder is comprised of a spatial decoder and a temporal decoder. Considering the dependency between duration and location, we expect to predict the location chain first and then utilize it to enhance the prediction of the duration chain. Inspired by the natural language process tasks, we add the start token $\rm<SOS>$ and end token $\rm<EOS>$ into the location chain to address the challenges of the unknown start location and the uncertain chain length. Besides, we employ the Transformer decoder [60] as the backbone of the spatial decoder, along with a positional encoding module and an FNN. The self-attention mechanism in the Transformer decoder and positional encodings enable the model to effectively capture long-range dependencies and maintain the temporal order of the input sequence, thereby alleviating accumulative errors in multi-step prediction tasks. Commencing from the start token, the spatial decoder utilizes the historical information and predicted results to make iterative predictions for the daily location chain. This process is formulated as:
@@
$\begin{aligned} \hat{\rm C}^{\rm dur}=\{\hat{t}_{1}\dots,\hat{t}_{m}\}={\rm TemporalDec}(\{\hat{s^{\prime}}_{1}\oplus z^{\rm dur},\dots,\hat{s^{\prime}}_{m}\oplus z^{\rm dur}\}),\end{aligned}$
@@ V-A1 Data Description
Experiments are conducted on four private  [^1] The dataset is not publicly disclosed due to privacy constraints. urban-scale datasets, Boston, Los Angeles (LA), San Francisco Bay Area (SFBay), Shanghai, collected from real-world mobile records, and one public dataset Tokyo [61], comprising synthetic mobility data. The preprocessing steps for these four CDR datasets follow [28]. Besides, we partition each city into uniform grids with grid sizes of 1 kilometer. As described in Preliminary, a day is divided into multiple time windows, and only the record with the longest stay duration is retained within each window. Boston, LA, and SFBay utilize half-hour windows, whereas Shanghai and Tokyo use hour-long windows.
The Tokyo dataset, as distinct from the other four datasets, undergoes two specific preprocessing steps. First, we identified users’ home locations by leveraging the approach in [28], retaining only those users whose home location could be determined, thus focusing on the resident population. A location is classified as “home” if it accounts for more than 30% of visits during nighttime hours (10PM–6AM). Second, we applied mobility behavior constraints inspired by [57, 28], ensuring that 90% of a user’s daily trajectories include fewer than 10 unique locations.
The statistics of the datasets are presented in Table I, where “# Days” means the average number of days and “# Admins” means the number of the administrative division. Boston, LA, and SFBay employ census tracts as administrative units, whereas Shanghai uses towns. In the Tokyo dataset, due to blurred exact coordinates, we partitioned administrative regions using 7x7 grid squares. The training, validation, and test sets are obtained by partitioning the original dataset in a ratio of 6:1:3 based on the time span.
@@
Boston LA SFBay Shanghai Tokyo
# Individuals 50,000 50,000 50,000 20,000 16,151
# Locations 10,294 6,188 8,032 16,050 7,153
# Admins 1,126 2,493 1,406 234 225
# Records 6,749,540 7,768,373 7,700,986 8,209,910 7,502,839
# Days 39 41 41 99 75
Time range
Feb. 20, 2010
to
Mar. 30, 2010
Oct. 15, 2012
to
Nov. 24, 2012
Oct. 15, 2012
to
Nov. 24, 2012
Jan. 1, 2014
to
Apr. 29, 2014
Unknown
@@ V-A2 Baselines
We compared 11 baseline models, which can be categorized into three groups based on their task orientation: next-location prediction models, successive location prediction models, and mid-term mobility prediction models. A brief overview of these baseline models is provided below.
Next-location prediction models
@@
Successive prediction model
@@
Mid-term mobility prediction model
  • HTAED [56] introduces an LSTM-based encoder-decoder model with a hierarchical temporal attention module for multi-day mobility prediction.
@@ V-A4 Implementation
These baselines are implemented with their open-source codes. Hyperparameter tuning is performed on the validation set to attain optimal performance. The implementation of MSTDP is based on the PyTorch framework. The heterogeneous GNNs adopt a two-layer structure, while the layer numbers of Transformer encoder and decoder are also set to 2. The embedding dimensions of location and time are 512 and 512. The hidden dimensions of location and time are 1024 and 512. The head numbers of attention in the Transformer encoder and decoder are set to 8. We apply the Adam optimizer to train MSTDP with a learning rate of $10^{-4}$. The hyperparameter $\lambda$ is set to 1. The number of training epochs is 60. Our source code is available at https://github.com/urbanmobility/MSTDP.git
@@ V-B Metric Comparison
We initially evaluate all models using datasets Boston, Shanghai, and Tokyo, and subsequently select the more recent baselines to compare against datasets LA and SFBay. The results are exhibited in Table II and Table III. From these results, we draw the following observations and analysis: (1) MSTDP demonstrates superior performance on both datasets, validating its effectiveness in mid-term mobility prediction tasks. Compared to the top-performing baseline, MSTDP exhibits improvements of 3.6% in Acc, 9.0% in DevDist, 20.3% in TravelDist and 19.4% in DepartTime. A smaller DevDist of MSTDP indicates closer proximity between the predicted and the actual locations. Furthermore, lower metrics for TravelDist and DepartTime demonstrate improved individual consistency between predicted and actual trajectories regarding travel time and distance patterns in our model. (2) Among these baselines, HTAED excels due to its dual-layer attention mechanism tailored for successive prediction, effectively weighing intra-day and inter-day behavioral patterns. In contrast, RLMob, which employs a reinforcement learning framework for successive prediction, falls short by relying solely on RNNs to model transition relationships, resulting in less favorable performances. These baselines, initially developed for next-location prediction, overlook temporal patterns such as daily and weekly cycles, which complicates their direct application to mid-term forecasting tasks. Moreover, when applied to Boston, LA, or SFBay datasets, these baseline predictions show significant performance degradation compared to Shanghai and Tokyo, notably highlighted by SNPM. This discrepancy primarily stems from varying time window sizes across datasets; for instance, Boston employs a half-hour window, leading to increased location redundancy and longer daily trajectories. These factors present substantial challenges for model performance, as discussed in Introduction I.
@@ V-C Statistical Physics Analysis
Origin-Destination (OD) flow pairs. We collect the grid-level OD flow pairs between the actual and predicted flows and illustrate them in Figure 7. We calculate $\rm{R}^{2}$ and CPC [64] between actual and predicted flows for numerical comparison. AGRAN and MTNet predict a limited number of flows, consistent with previous analyses indicating their preference for predicting full-day stays. HTAED overestimates flows with an $\rm R^{2}$ of -0.442 and $\rm CPC$ of 0.461, whereas our MSTDP model achieves closer predictions to actual flow quantities, with an improved $\rm R^{2}$ of 0.728 and $\rm CPC$ of 0.672.
@@
Origin-Destination flows in map. Figure 8 illustrates the actual and predicted OD flows in the map, where the trips between grids are aggregated at the census tract level. The global travel patterns predicted by the three baseline models deviate significantly from the actual results. HTAED achieves $\rm R^{2}$ and CPC values of 0.534 and 0.629, whereas MSTDP achieves 0.872 and 0.792, indicating performance improvements of 63.3% and 25.9% in our model. In this subsection, we examine OD flows at the grid and tract levels. Additionally, Figure 9 illustrates OD flows at the zip code level. MSTDP achieves R-squared ($R^{2}$) and CPC values of 0.969 and 0.877, while HTAED obtains 0.928 and 0.817. The performance improvements are 4.4% and 7.3%, respectively. This observation suggests that larger spatial aggregations may obscure differences between models.
@@ V-D Ablation Study
We compare MSTDP with three variants to examine the effectiveness of the modules in MSTDP. The ablated variants of MSTDP are as follows: (1) MSTDP-D: it eliminates the spatial-temporal decoupler and directly encodes the long trajectories; (2) MSTDP-H: it abandons the hierarchical structure and solely utilizes the daily encoder; (3) MSTDP-G: it replaces the graph-based location representation with a word embedder; (4) MSTDP-LD: the temporal decoder excludes the concatenation from the location decoder and uses a Transformer decoder structure rather than the Transformer encoder.
@@
Next Day Prediction
MSTDP-D 0.571 3.891 0.539 0.560
@@
Next Week Prediction
MSTDP-D 0.404 5.634 0.564 0.570
@@ V-E Hyperparameter Analysis
To examine MSTDP’s sensitivity to hyperparameters, we vary four hyperparameters to observe performance fluctuations: $\lambda$ (the weight of the temporal loss), head number (#Head\#Head) in the Transformer encoder, embedding dimensions (Dim), and layer depths (#Layer\#Layer). For $\lambda$, we test values of 0.01, 0.1, 1, 10, and 100, which impact the optimization balance between spatial and temporal losses. For head number, we evaluate 2, 4, and 8 attention heads. Additionally, we investigate the effect of embedding dimensions, testing values of 256, 512, and 768, and layer depths, comparing configurations with 1, 2, and 3 layers. The default parameter settings are $\lambda=1$, #Head=8\#Head=8, $Dim=512$, and #Layer=2\#Layer=2, with training epochs at 40.
Table V presents the sensitivity analysis of MSTDP under various hyperparameter configurations, revealing several key trends. For the temporal loss weight ($\lambda$), larger values consistently yield better results. $\lambda=10$ achieves the highest Acc, while $\lambda=100$ improves TravelTime. Smaller values, such as $\lambda=0.01$ and $\lambda=0.1$, lead to lower accuracy and higher deviation distance, emphasizing the importance of effectively balancing the temporal loss. Increasing the attention head number (#Head\#Head) improves performance, with the default setting (#Head=8\#Head=8) achieving the lowest DevDist and TravelDist, demonstrating the advantage of richer attention mechanisms. For layer depth (#Layer\#Layer) and embedding dimensions ($Dim$), the default configurations of 2 layers and $Dim=512$ achieve the best overall balance. While deeper or wider models might benefit from extended training (more epochs), they also pose a heightened risk of overfitting. The default configuration achieves optimal performance across all metrics, highlighting the importance of a carefully tuned model for mid-term mobility prediction tasks.
@@
Next Day Prediction
@@
#Head=2\#Head=2 0.568 3.826 0.392 0.426
#Head=4\#Head=4 0.571 3.803 0.389 0.430
#Layer=1\#Layer=1 0.562 3.875 0.395 0.428
#Layer=3\#Layer=3 0.574 3.796 0.396 0.435
$Dim=256$ 0.565 3.891 0.399 0.421
$Dim=768$ 0.567 3.859 0.416 0.444
Default 0.574 3.760 0.383 0.424
@@ V-F Application to Epidemic Transmission
The SEIR model is commonly used to represent the epidemic spreading in a population [65, 26]. It categorizes all individuals into four states: susceptible (S), exposed (E), infectious (I), and removed (R). Susceptible individuals can become infected through close contact with those carrying the infection and subsequently transition into the exposed state. There exists a probability of transitioning from the exposed state to infection. Lastly, infectious individuals transition into the removed state with a probability. The removed state assumes that individuals are no longer infected and cannot transmit the infection to others. Specifically, the probability of being exposed is formulated as $\rm P_{exp,i}=\frac{\alpha\beta I_{i}}{\rm N_{i}}$, where $\alpha$ is the number of close contacts, $\beta$ is the infectious rate, $\rm I_{i}$ and $\rm N_{i}$ stand for the number of infectious people and total population in the $i$-th subregion of the city. The infectious probability $\rm P_{inf}$ is the inverse of the average incubation period, while the removed probability $\rm P_{rem}$ is the inverse of the infectious period. The detailed parameter settings are illustrated in Table VI. $\rm R_{0}$, the basic reproduction number, represents the average number of secondary infections caused by an infectious individual. It is calculated as the product of $\alpha$$\beta$, and the average infectious period.
@@
In this study, due to the sparse population in the dataset, we initially aggregate grid-level trajectories to the admin level (census tract for Boston) and assign an SEIR epidemic model to each administrative area. Subsequently, we employ these trajectories to construct a population transition probability matrix for each time window. These matrices are normalized to represent the probability of individuals choosing to stay or move between administrative areas during different periods. The population of each administrative area is obtained from Worldpop  [^2] https://www.worldpop.org/. For each time window, we first employ the SEIR model to calculate the status change of individuals within each administrative area, and then determine individual movement based on the transition probability matrix. All individuals are considered independent of each other.
@@
V-G Training and Inference Time Comparison
To evaluate the practical efficiency of these models, we compared the training and inference times on the Boston dataset. The experiments were conducted on an NVIDIA 4090 GPU, with the model configurations kept consistent with those used in the performance comparison experiments. Figure 12 presents the training time results, while Table VII summarizes the inference time comparison. For inference time, we measured the duration required to predict daily mobility for 1,000 users with a batch size of 50. Results demonstrate that HTAED achieves the fastest training and inference times, owing to its lightweight model structure. MSTDP significantly reduces training time by 79.4% and 44.6% compared to MTNet and AGRAN, respectively. Similarly, MSTDP achieves impressive inference time reductions of 99.5% and 99.8% compared to MTNet and AGRAN. These findings highlight MSTDP’s ability to balance computational efficiency with high predictive performance.
Figure 12: Comparison of loss reduction across epochs on Boston dataset. The loss values are min-max normalized for better visualization.
TABLE VII: Comparison of inference time on Boston dataset.
MSTDP HTEAD MTNet AGRAN
Infer time (s) 10.1 1.4 1956.0 6853.7
@@ VI Conclusion
While our study has made significant progress in addressing mid-term mobility prediction, there remain several important directions for future exploration. First, MSTDP achieves strong performance on regular movement patterns and dense datasets; however, its generalization to irregular movement scenarios or sparse data environments presents opportunities for further improvement. Investigating adaptive modeling strategies, such as dynamic learning mechanisms or the incorporation of additional contextual factors (e.g., environmental or social information), could enhance the versatility and robustness of the model in such conditions. Second, the use of mobile phone data in trajectory prediction inherently involves sensitive user information, raising potential privacy concerns. Although all datasets in this study were anonymized, future work could explore integrating advanced privacy-preserving mechanisms to ensure robust data security during deployment in real-world applications. Third, real-world mobility behaviors are shaped by various external influences, such as weather patterns, public health emergencies, holidays, and major events. Future efforts could integrate these external data sources and develop dynamic mechanisms to adapt predictions based on such factors, thereby improving the model’s applicability in diverse and evolving scenarios. Finally, an important aspect of mid-term mobility prediction is the challenge of cumulative error propagation in multi-step predictions, which impacts accuracy over longer time horizons. Future research could focus on designing methodologies to further enhance the accuracy and reliability of multi-day trajectory forecasting.